OpenStudy (anonymous):

trying to find the x-intercepts of y=-(x-1)^3 -1

OpenStudy (anonymous):

To find x-intercepts, set y = 0. -(x-1)^3 - 1 = 0 Add 1 to both sides, you get -(x-1)^3 = 1 Divide both sides by -1, you get (x - 1)^3 = -1 What number cubed is -1? Answer is -1, becuase (-1)^3 = -1 so set x - 1 = -1 x = 0 Therefore, x = 0 is the ONLY x-intercept.

OpenStudy (anonymous):

wow. For some reason I don't know why I could not set that up correctly. Thank you

OpenStudy (anonymous):

Welcome.